# On the Weierstraß form of infinite dimensional differential algebraic equations

### v1

@@ Abstract
The solvability for infinite dimensional differential algebraic equations possessing a resolvent index and a Weierstraß form is studied. In particular, the concept of integrated semigroups is used to determine a subset on which solutions exist and are unique. This information is later used for a important class of systems, namely port-Hamiltonian differential algebraic equations.
@@ 1 Introduction
where $J$ is in Jordan form and $N$ is nilpotent, as long as $(\lambda E-A)^{-1}\in L(\mathcal{Z},\mathcal{X})$ for some $\lambda\in\mathbb{C}$, [10, Ch. 2.1]. In this context, the nilpotency degree of $N$ is called the nilpotency index of $(E,A)$ [2, 10]. For finite-dimensional spaces, this definition of index coincides with other common definitions of the index. In this context, investigation of the different index definitions can be found in [4, 21, 22] for the resolvent index, [6, 18] for the radiality index, [2] for a general comparison of different known indices in infinite-dimensions and [10, 11] for the finite-dimensional case.
@@
for some $C,\omega>0$ [14, Thm. 4.8]. For $n=0$ this matches the Hille-Yosida resolvent condition used for the generation of a $C_{0}$-semigroup. Notice, that a $C_{0}$-semigroup is nothing else than a 00-times integrated semigroup and (3) coincides with the complex resolvent index for $E=I$. Based on this, we build on existing solution methods for the abstract Cauchy problem $\frac{d}{dt}x(t)=Ax(t)$, $x(0)=x_{0}$ on the subspace $\mathcal{D}(A^{n})={\rm ran}\,((\lambda I-A)^{-1})^{n}$ [9, §1.3] and extend these studies to the differential algebraic equation (1) with $x_{0}\in{\rm ran}\,((\lambda E-A)^{-1}E)^{n}$. For further details on integrated semigroups we refer to [1, 14].
@@
The paper is organized as follows. In Section 2 a set of initial conditions for which (1) has a solution is characterized. In Section 3 the radiality index is defined and it is shown that if if ${\rm ran}\,E$ is closed, existence of this index implies existence of a Weierstraß form. Several new results relating the radiality index and the degree of nilpotency are also obtained. Furthermore, provided that the complex resolvent index exists, and the Weierstraß form exists with densely defined $A_{1}$, then $A_{1}$ generates an integrated semigroup.
@@ 2 Existence of solutions on a subspace
In this context, we say that $x\colon\mathbb{R}_{\geq 0}\to\mathcal{X}$ is a classical solution of (1) for an initial condition $x_{0}\in\mathcal{X}$, if $Ex(\cdot)$ is continuously differentiable on $\mathbb{R}_{>0}$, $x(t)\in\mathcal{D}(A)$ for all $t\geq 0$, $x(0)=x_{0}$ and (1) holds, and we say that $x\colon\mathbb{R}_{\geq 0}\to\mathcal{X}$ is a mild solution, if $Ex(\cdot)$ is continuous, and for almost all $t\geq 0$, it holds $\int_{0}^{t}x(s)\,\mathrm{d}s\in\mathcal{D}(A)$ with
@@
The following theorem is a generalization of the solution theory of the abstract Cauchy problem $\frac{d}{dt}x(t)=Ax(t)$, $x(0)=x_{0}$ and extends the proofs of [9, p. 34] and [15, Ch. 4, Thm. 1.2] to the DAE case with the help of pseudo-resolvents.
@@
Assume, that $(E,A)$ has a complex resolvent index. Then, for every $x_{0}\in{\rm ran}\,((\mu E-A)^{-1}E)^{p_{\mathrm{c,res}}^{(E,A)}+2}$ there exists a unique solution $x\colon\mathbb{R}_{\geq 0}\to\mathcal{X}$ of (1).
@@
Let $x_{0}\in{\rm ran}\,R(\mu)^{p}$. Thus, there exists a $z_{0}\in\mathcal{X}$ and a $\mu\in\mathbb{C}_{{\rm Re}>\omega}$ such that $x_{0}=(-1)^{p-1}R(\mu)^{p}z_{0}$. Hence,
$R(\lambda)x_{0}=\frac{R(\lambda)z_{0}}{(\lambda-\mu)^{p}}-\frac{R(\mu)z_{0}}{(\lambda-\mu)^{p}}-\ldots-\frac{R(\mu)^{p}z_{0}}{(\lambda-\mu)}=\frac{R(\lambda)z_{0}}{(\lambda-\mu)^{p}}-\sum_{k=1}^{p}\frac{R(\mu)^{k}z_{0}}{(\lambda-\mu)^{p-k+1}}$ (8)
for all $\lambda\in\rho(E,A)$ with $\lambda\neq\mu$. This holds, since $R(\mu)$ is a pseudo-resolvent and, in particular, fulfils the resolvent identity. We want to multiply both sides of (8) with $\mathrm{e}^{\lambda t}$ and integrate along the imaginary axis at ${\rm Re}\lambda=\omega$. Now, for a fixed $k\in\{1,\ldots,p\}$ the terms $\frac{R(\mu)^{k}}{(\lambda-\mu)^{p-k+1}}$ converge to 00 uniformly for $\lambda\to\infty$. Thus, by Jordan’s Lemma, the following integrals vanish
$\int_{\omega-i\infty}^{\omega+i\infty}\mathrm{e}^{\lambda t}\frac{R(\mu)^{k}z_{0}}{(\lambda-\mu)^{p-k+1}}\,\mathrm{d}\lambda,$
for $k\in\{1,\ldots,p\}$. By (7) we conclude
@@
for a $\tilde{C}>0$ and for all $\lambda\neq\mu$. Thus, the integral of $\mathrm{e}^{\lambda t}\frac{R(\lambda)z_{0}}{(\lambda-\mu)^{p}}$ exists and, especially, the function
$x(t)\coloneqq-\frac{1}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\mathrm{e}^{\lambda t}R(\lambda)x_{0}\,\mathrm{d}\lambda=-\frac{1}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\mathrm{e}^{\lambda t}\frac{R(\lambda)z_{0}}{(\lambda-\mu)^{p}}\,\mathrm{d}\lambda,\quad t\geq 0,$ (10)
@@
$\frac{d}{dt}Ex(t)$ $=-\frac{1}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\lambda\mathrm{e}^{\lambda t}\frac{ER(\lambda)z_{0}}{(\lambda-\mu)^{p}}\,\mathrm{d}\lambda$
$=-\frac{1}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\mathrm{e}^{\lambda t}\frac{AR(\lambda)z_{0}}{(\lambda-\mu)^{p}}\,\mathrm{d}\lambda-\underbrace{\frac{1}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\mathrm{e}^{\lambda t}\frac{Ez_{0}}{(\lambda-\mu)^{p}}\,\mathrm{d}\lambda}_{=0}=Ax(t).$
Furthermore,
$x(0)=-\frac{1}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\frac{R(\lambda)z_{0}}{(\lambda-\mu)^{p}}\,\mathrm{d}\lambda.$
@@
$x(0)=\frac{1}{(p-1)!}\lim_{\lambda\to\mu}\frac{d^{p-1}}{d\lambda^{p-1}}R(\lambda)z_{0}=(-1)^{p-1}R(\mu)^{p}z_{0}=x_{0}.$
Here, we used $\frac{d}{d\lambda}R(\lambda)^{n}z_{0}=-nR(\lambda)^{n+1}z_{0}$ for all $n\in\mathbb{N}$ and $\lambda\in\rho(E,A)$ ([18, Lem. 2.4.1]). Thus, $x$ is a classical solution of (1).
Uniqueness: We assume that $x$ is a solution with $x(0)=0$ and prove that $x(0)=0$ for all $t\geq 0$. Since $x$ is a solution of (1)
$\frac{d}{dt}R(\lambda)x(t)=(\lambda E-A)^{-1}Ax(t)=R(\lambda)x(t)-x(t),$
@@
$\lim_{{\rm Re}\lambda\to\infty}\bigg{\|}\int_{0}^{t-\sigma}$ $\mathrm{e}^{\lambda(t-\sigma-\tau)}x(t)\,\mathrm{d}\tau\bigg{\|}$
$=\lim_{{\rm Re}\lambda\to\infty}\mathrm{e}^{-\sigma{\rm Re}\lambda}\left\|\int_{0}^{t-\sigma}\mathrm{e}^{\lambda(t-\tau)}x(t)\,\mathrm{d}\tau\right\|$
$\leq\lim_{{\rm Re}\lambda\to\infty}\mathrm{e}^{-\sigma{\rm Re}\lambda}\bigg{\|}\underbrace{\int_{0}^{t}\mathrm{e}^{\lambda(t-\tau)}x(t)\,\mathrm{d}\tau}_{=R(\lambda)x(t)}\bigg{\|}+\underbrace{\mathrm{e}^{-\sigma{\rm Re}\lambda}\left\|\int_{t-\sigma}^{t}\mathrm{e}^{\lambda(t-\tau)}x(t)\,\mathrm{d}\tau\right\|}_{\to 0}=0.$
@@
and by [15, Ch. 4, Lem. 1.1] $x(\tau)=0$ for all $0\leq\tau\leq t-\sigma$. Since $\sigma$ and $t$ were arbitrary, we deduce $x(\tau)=0$ for all $\tau\geq 0$. ∎
@@
    $=\int_{0}^{t}\bigg{(}-\frac{1}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\lambda\mathrm{e}^{\lambda s}\frac{R(\lambda)z_{0}}{(\lambda-\mu)^{p}}\,\mathrm{d}\lambda\bigg{)}\,\mathrm{d}s$
@@
    was analyzed in more detail. To be more precise, it was shown that $E^{-1}A$ generates a $C_{0}$-semigroup on $\overline{{\rm ran}\,((\mu E-A)^{-1}E)^{p_{\rm c,res}^{(E,A)}+2}}$ if and only if $U$ is closed and for each $x_{0}\in U$ the mild solution is unique.The assumption that $U$ is closed is restrictive and it also requires information about $U$.
@@
    In [4, Section 8] the existence of solutions on the subspace $\overline{{\rm ran}\,R(\lambda)^{k}}$, for $R(\lambda)\coloneqq(\lambda E-A)^{-1}E$, has been studied under slightly stronger conditions compared to the ones in Theorem 2.2. More specifically, under the assumption that there exists $k\in\mathbb{N}$ with $k\geq 1$, $\omega\in\mathbb{R}$ and $M>0$, such that $(\omega,\infty)\subseteq\rho(E,A)$ and
@@ 3 Sufficient condition for a Weierstraß form
We call two differential-algebraic equations $\frac{d}{dt}Ex=Ax$ and $\frac{d}{dt}\tilde{E}\tilde{x}=\tilde{A}\tilde{x}$ defined on Hilbert spaces $\mathcal{X}$,$\mathcal{Z}$ and $\tilde{X}$, $\tilde{\mathcal{Z}}$ equivalent, denoted by $(E,A)\sim(\tilde{E},\tilde{A})$, if there are two bounded isomorphisms $P\colon\mathcal{X}\to\tilde{X}$, $Q\colon\mathcal{Z}\to\tilde{Z}$, such that $E=Q^{-1}\tilde{E}P$ and $A=Q^{-1}\tilde{A}P$. Based on that, we define the notion of a Weierstraß form as follows.
@@
where $N\colon\mathcal{Y}^{2}\to\mathcal{Y}^{2}$ is a nilpotent operator, $A_{1}\colon\mathcal{D}(A_{1})\subseteq\mathcal{Y}^{1}\to\mathcal{Y}^{1}$ is a linear operator and $I_{\mathcal{Y}^{i}}$ indicated the identity operator on the associated subspace $\mathcal{Y}^{i}$, $i=1,2$. Furthermore, the nilpotency degree of $N$ is known as the nilpotency index $p_{\rm nilp}^{(E,A)}$ [2, Section 2].
@@
By [18, Section 2.1, 2.2 & 2.5] these spaces are independent of the choice of $\lambda_{i}$, $i\in\mathopen{}\left\{0,\ldots,p\right\}\mathclose{}$,
@@
is closed and densely defined, whereby $\mathcal{D}(A_{i})\coloneqq\mathcal{D}(A)\cap\mathcal{Z}^{i}$, $i\in\mathopen{}\left\{1,2\right\}\mathclose{}$. This was already proved for $p=0$ in [6, Prop. 2.5] and under stronger radiality assumptions in [18, Cor. 2.5.2, Thm. 2.5.3]. Let $x\in\mathcal{X}^{1}$. Since $\mathcal{X}=\mathcal{X}^{1}\oplus\mathcal{X}^{2}$, $x\in\mathcal{X}^{1}$ if and only if $Px=x$. By (17) we conclude
@@
What is left to show is, that $E_{1}$ is boundedly invertible. Since $E_{1}$ is bounded, ${\rm ker}\,E_{1}\subseteq{\rm ker}\,E\subseteq\mathcal{X}^{2}$ and $\mathcal{X}=\mathcal{X}^{1}\oplus\mathcal{X}^{2}$, it follows that $E_{1}$ is injective. What remains to show is the surjectivity of $E$, because then the assertion follows from the Closed Graph Theorem. Since the space $\mathcal{Z}^{1}$ does not depend on the choice of the $\lambda_{i}$, we can choose a $\lambda\in\rho(E,A)$, such that $\mathcal{Z}^{1}=\overline{{\rm ran}\,(E(\lambda E-A)^{-1})^{p+2}}^{\|\cdot\|_{\mathcal{Z}}}$ [18, Lem. 2.2.8]. Let $y\in{\rm ran}\,(E(\lambda E-A)^{-1})^{p+2}$. Then there exists a $z\in\mathcal{Z}$ with
$y=(E(\lambda E-A)^{-1})^{p+2}z=E((\lambda E-A)^{-1}E)^{p+1}(\lambda E-A)^{-1}z.$
Given that $(\lambda E-A)^{-1}z\in\mathcal{D}(A)$ we have
@@
$\frac{d}{dt}Ex(t)=Ax(t)$$\Longleftrightarrow$ $\frac{d}{dt}\tilde{R}E\tilde{P}^{-1}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$ $=\tilde{R}A\tilde{P}^{-1}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$ (18e)
$\Longleftrightarrow$ $\frac{d}{dt}\begin{bmatrix}E_{1}&0\\ 0&E_{2}\end{bmatrix}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$ $=\begin{bmatrix}A_{1}&0\\ 0&A_{2}\end{bmatrix}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$ (18n)
$\Longleftrightarrow$ $\frac{d}{dt}\begin{bmatrix}I_{\mathcal{Z}^{1}}&0\\ 0&E_{2}A_{2}^{-1}\end{bmatrix}\begin{bmatrix}z_{1}(t)\\ z_{2}(t)\end{bmatrix}$ $=\begin{bmatrix}A_{1}E_{1}^{-1}&0\\ 0&I_{\mathcal{Z}^{2}}\end{bmatrix}\begin{bmatrix}z_{1}(t)\\ z_{2}(t)\end{bmatrix},$ (18w)
where $\tilde{P}x(t)=\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$ and $\begin{bmatrix}z_{1}(t)\\ z_{2}(t)\end{bmatrix}=\begin{bmatrix}E_{1}^{-1}x_{1}(t)\\ A_{2}^{-1}x_{2}(t)\end{bmatrix}$. The uniqueness of the nilpotency index follows from [2, Section 6]
@@
If we consider $E\frac{d}{dt}x=Ax$ instead of $\frac{d}{dt}Ex=Ax$ in the proof of Theorem 3.3, then
$E\frac{d}{dt}x(t)=Ax(t)$$\Longleftrightarrow$ $\tilde{R}E\tilde{P}^{-1}\frac{d}{dt}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$ $=\tilde{R}A\tilde{P}^{-1}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$ (19e)
$\Longleftrightarrow$ $\begin{bmatrix}E_{1}&0\\ 0&E_{2}\end{bmatrix}\frac{d}{dt}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$ $=\begin{bmatrix}A_{1}&0\\ 0&A_{2}\end{bmatrix}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$ (19n)
$\Longleftrightarrow$ $\begin{bmatrix}I_{\mathcal{X}^{1}}&0\\ 0&A_{2}^{-1}E_{2}\end{bmatrix}\frac{d}{dt}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$ $=\begin{bmatrix}E_{1}^{-1}A_{1}&0\\ 0&I_{\mathcal{X}^{2}}\end{bmatrix}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix},$ (19w)
where $\tilde{P}x(t)=\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$.
@@
Let $\mathcal{W}$ be a Hilbert space and $(A_{0},\mathcal{D}(A_{0}))$ generate a $C_{0}$-semigroup on $\mathcal{W}$. For $b\in\mathcal{D}(A)$ and $c\in\mathcal{D}(A_{0}^{\ast})$ with $\langle b,c\rangle\neq 0$ define the operators $B_{u}=bu$, where $u\in\mathbb{C}$ and $Cz=\langle z,c\rangle$ for any $z\in\mathcal{W}$. We define the following DAE on $\mathcal{Z}=\mathcal{W}\times\mathbb{C}$ by

### v2

@@ Abstract
The solvability for infinite dimensional differential algebraic equations possessing a resolvent index and a Weierstraß form is studied. In particular, the concept of integrated semigroups is used to determine a subset on which solutions exist and are unique. This information is later used for a important class of systems, namely, port-Hamiltonian differential algebraic equations.
@@ 1 Introduction
where $J$ is in Jordan form and $N$ is nilpotent, as long as $(\lambda E-A)^{-1}\in L(\mathcal{Z},\mathcal{X})$ for some $\lambda\in\mathbb{C}$, [11, Ch. 2.1]. In this context, the nilpotency degree of $N$ is called the nilpotency index of $(E,A)$ [2, 11]. For finite-dimensional spaces, this definition of index coincides with other common definitions of the index. In this context, investigation of the different index definitions can be found in [5, 22, 23] for the resolvent index, [7, 19] for the radiality index, [2] for a general comparison of different known indices in infinite-dimensions and [11, 12] for the finite-dimensional case.
@@
for some $C,\omega>0$ [15, Thm. 4.8]. For $n=0$ this matches the condition of a sectorial operator operator $A$, which is used for the generation of a analytic semigroup. Notice, that (3) coincides with the complex resolvent index for $E=I$. Based on this, we build on existing solution methods for the abstract Cauchy problem $\frac{d}{dt}x(t)=Ax(t)$, $x(0)=x_{0}$ on the subspace $\mathcal{D}(A^{n})={\rm ran}\,((\lambda I-A)^{-1})^{n}$ [10, §1.3] and extend these studies to the differential algebraic equation (1) with $x_{0}\in{\rm ran}\,((\lambda E-A)^{-1}E)^{n}$. For further details on integrated semigroups we refer to [1, 15].
@@
The paper is organized as follows. In Section 2 a set of initial conditions for which (1) has a solution is characterized. In Section 3 the radiality index is defined and it is shown that if ${\rm ran}\,E$ is closed, existence of this index implies existence of a Weierstraß form. Several new results relating the radiality index and the degree of nilpotency are also obtained. Furthermore, provided that the complex resolvent index exists, and the Weierstraß form exists with densely defined $A_{1}$, then $A_{1}$ generates an integrated semigroup.
@@ 2 Existence of solutions on a subspace
In this context, we say that $x\colon\mathbb{R}_{\geq 0}\to\mathcal{X}$ is a classical solution of (1) for an initial condition $x_{0}\in\mathcal{X}$, if $x(\cdot)$ is continuous on $\mathbb{R}_{\geq 0}$, $Ex(\cdot)$ is continuously differentiable on $\mathbb{R}_{>0}$, $x(t)\in\mathcal{D}(A)$ for all $t\geq 0$, $x(0)=x_{0}$ and (1) holds, and we say that $x\colon\mathbb{R}_{\geq 0}\to\mathcal{X}$ is a mild solution, if $Ex(\cdot)$ is continuous, and for almost all $t\geq 0$, it holds $\int_{0}^{t}x(s)\,\mathrm{d}s\in\mathcal{D}(A)$ with
@@
The following theorem is a generalization of the solution theory of the abstract Cauchy problem $\frac{d}{dt}x(t)=Ax(t)$, $x(0)=x_{0}$ and extends the proofs of [10, p. 34] and [16, Ch. 4, Thm. 1.2] to the DAE case with the help of pseudo-resolvents.
@@
Assume, that $(E,A)$ has a complex resolvent index. Then, for every $x_{0}\in{\rm ran}\,((\mu E-A)^{-1}E)^{p_{\mathrm{c,res}}^{(E,A)}+2}$ there exists a unique classical solution $x\colon\mathbb{R}_{\geq 0}\to\mathcal{X}$ of (1).
@@
Let $x_{0}\in{\rm ran}\,R(\mu)^{p}$. Thus, there exists a $z_{0}\in\mathcal{X}$ and a $\mu\in\mathbb{C}_{{\rm Re}>\omega}$ such that $x_{0}=(-1)^{p-1}R(\mu)^{p}z_{0}$ and therefore
@@
for a $\tilde{C}>0$ and for all $\lambda\neq\mu$. Thus, the integral of $\mathrm{e}^{\lambda t}\frac{R(\lambda)z_{0}}{(\lambda-\mu)^{p}}$ along the imaginary axis at ${\rm Re}\,\lambda=\omega$ exists and, especially, the function
$x(t)\coloneqq\frac{(-1)^{p}}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\mathrm{e}^{\lambda t}\frac{R(\lambda)z_{0}}{(\lambda-\mu)^{p}}\,\mathrm{d}\lambda,\quad t\geq 0,$ (9)
@@
$\frac{d}{dt}Ex(t)$ $=\frac{(-1)^{p}}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\lambda\mathrm{e}^{\lambda t}\frac{ER(\lambda)z_{0}}{(\lambda-\mu)^{p}}\,\mathrm{d}\lambda$ (10)
$=\frac{(-1)^{p}}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\mathrm{e}^{\lambda t}\frac{AR(\lambda)z_{0}}{(\lambda-\mu)^{p}}\,\mathrm{d}\lambda+\underbrace{\frac{(-1)^{p}}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\mathrm{e}^{\lambda t}\frac{Ez_{0}}{(\lambda-\mu)^{p}}\,\mathrm{d}\lambda}_{=0}=Ax(t).$ (11)
Here, we used that $\frac{Ez_{0}}{(\lambda-\mu)^{p}}$ converges uniformly to 00 for $\lambda\to\infty$ and Jordan’s Lemma to show that the second integral in (11) vanishes. Furthermore,
$x(0)=\frac{(-1)^{p}}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\frac{R(\lambda)z_{0}}{(\lambda-\mu)^{p}}\,\mathrm{d}\lambda.$
@@
$x(0)=\frac{(-1)^{p-1}}{(p-1)!}\lim_{\lambda\to\mu}\frac{d^{p-1}}{d\lambda^{p-1}}R(\lambda)z_{0}=(-1)^{p-1}R(\mu)^{p}z_{0}=x_{0}.$
Here, we used $\frac{d}{d\lambda}R(\lambda)^{n}z_{0}=-nR(\lambda)^{n+1}z_{0}$ for all $n\in\mathbb{N}$ and $\lambda\in\rho(E,A)$ ([19, Lem. 2.4.1]). Thus, $x$ is a classical solution of (1).
Uniqueness: We assume that $x$ is a solution with $x(0)=0$ and prove that $x(t)=0$ for all $t\geq 0$. Since $x$ is a solution of (1), we deduce
$\frac{d}{dt}R(\lambda)x(t)=(\lambda E-A)^{-1}Ax(t)=\lambda R(\lambda)x(t)-x(t),$
@@
$\lim_{{\rm Re}\lambda\to\infty}\bigg{\|}\int_{0}^{t-\sigma}$ $\mathrm{e}^{\lambda(t-\sigma-\tau)}x(\tau)\,\mathrm{d}\tau\bigg{\|}$
$=\lim_{{\rm Re}\lambda\to\infty}\mathrm{e}^{-\sigma{\rm Re}\lambda}\left\|\int_{0}^{t-\sigma}\mathrm{e}^{\lambda(t-\tau)}x(\tau)\,\mathrm{d}\tau\right\|$
$\leq\lim_{{\rm Re}\lambda\to\infty}\mathrm{e}^{-\sigma{\rm Re}\lambda}\bigg{\|}\underbrace{\int_{0}^{t}\mathrm{e}^{\lambda(t-\tau)}x(\tau)\,\mathrm{d}\tau}_{=R(\lambda)x(t)}\bigg{\|}+\underbrace{\mathrm{e}^{-\sigma{\rm Re}\lambda}\left\|\int_{t-\sigma}^{t}\mathrm{e}^{\lambda(t-\tau)}x(\tau)\,\mathrm{d}\tau\right\|}_{\to 0}=0.$
@@
and by [16, Ch. 4, Lem. 1.1] $x(\tau)=0$ for all $0\leq\tau\leq t-\sigma$. Since $\sigma$ and $t$ were arbitrary, we deduce $x(\tau)=0$ for all $\tau\geq 0$. ∎
@@
    $=\int_{0}^{t}\bigg{(}-\frac{1}{2\pi i}\int_{\omega-i\infty}^{\omega+i\infty}\lambda\mathrm{e}^{\lambda s}\frac{ER(\lambda)z_{0}}{(\lambda-\mu)^{p}}\,\mathrm{d}\lambda\bigg{)}\,\mathrm{d}s$
@@
    was analyzed in more detail. To be more precise, it was shown that $E^{-1}A$ generates a $C_{0}$-semigroup on $\overline{{\rm ran}\,((\mu E-A)^{-1}E)^{p_{\rm c,res}^{(E,A)}+2}}$ if and only if $U$ is closed and for each $x_{0}\in U$ the mild solution is unique. The assumption that $U$ is closed is a bit restrictive and it also requires information about $U$.
@@
    In [5, Section 8] the existence of solutions on the subspace $\overline{{\rm ran}\,R(\lambda)^{k}}$, for $R(\lambda)\coloneqq(\lambda E-A)^{-1}E$, has been studied under slightly stronger conditions compared to the ones in Theorem 2.2. More specifically, under the assumption that there exists $k\in\mathbb{N}$ with $k\geq 1$, $\omega\in\mathbb{R}$ and $M>0$, such that $(\omega,\infty)\subseteq\rho(E,A)$ and
@@ 3 Sufficient condition for a Weierstraß form
We call two differential-algebraic equations $\frac{d}{dt}Ex=Ax$ and $\frac{d}{dt}\tilde{E}\tilde{x}=\tilde{A}\tilde{x}$ defined on Hilbert spaces $\mathcal{X}$,$\mathcal{Z}$ and $\tilde{\mathcal{X}}$, $\tilde{\mathcal{Z}}$ equivalent, denoted by $(E,A)\sim(\tilde{E},\tilde{A})$, if there are two bounded isomorphisms $P\colon\mathcal{X}\to\tilde{\mathcal{X}}$, $Q\colon\mathcal{Z}\to\tilde{\mathcal{Z}}$, such that $E=Q^{-1}\tilde{E}P$ and $A=Q^{-1}\tilde{A}P$. Based on that, we define the notion of a Weierstraß form as follows.
@@
where $N\colon\mathcal{Y}^{2}\to\mathcal{Y}^{2}$ is a nilpotent operator, $A_{1}\colon\mathcal{D}(A_{1})\subseteq\mathcal{Y}^{1}\to\mathcal{Y}^{1}$ is a linear operator and $I_{\mathcal{Y}^{i}}$ indicates the identity operator on the associated subspace $\mathcal{Y}^{i}$, $i=1,2$. Furthermore, the nilpotency degree of $N$ is known as the nilpotency index $p_{\rm nilp}^{(E,A)}$ [2, Section 2].
@@
By [19, Section 2.1, 2.2 & 2.5] these spaces are independent of the choice of $\lambda_{i}$, $i\in\mathopen{}\left\{0,\ldots,p\right\}\mathclose{}$,
@@
is closed and densely defined, whereby $\mathcal{D}(A_{i})\coloneqq\mathcal{D}(A)\cap\mathcal{Z}^{i}$, $i\in\mathopen{}\left\{1,2\right\}\mathclose{}$. This was already proved for $p=0$ in [7, Prop. 2.5] and under stronger radiality assumptions in [19, Cor. 2.5.2, Thm. 2.5.3]. Let $x\in\mathcal{X}^{1}$. Since $\mathcal{X}=\mathcal{X}^{1}\oplus\mathcal{X}^{2}$, $x\in\mathcal{X}^{1}$ if and only if $Px=x$. By (18) we conclude
@@
What is left to show is, that $E_{1}$ is boundedly invertible. Since $E_{1}$ is bounded, ${\rm ker}\,E_{1}\subseteq{\rm ker}\,E\subseteq\mathcal{X}^{2}$ and $\mathcal{X}=\mathcal{X}^{1}\oplus\mathcal{X}^{2}$, it follows that $E_{1}$ is injective. What remains to show is the surjectivity of $E$, because then the assertion follows from the Closed Graph Theorem. Since the space $\mathcal{Z}^{1}$ does not depend on the choice of the $\lambda_{i}$, we can choose a $\lambda\in\rho(E,A)$, such that $\mathcal{Z}^{1}=\overline{{\rm ran}\,(E(\lambda E-A)^{-1})^{p+2}}^{\|\cdot\|_{\mathcal{Z}}}$ [19, Lem. 2.2.8]. Let $y\in{\rm ran}\,(E(\lambda E-A)^{-1})^{p+2}$. Then there exists a $z\in\mathcal{Z}$ with
$y=(E(\lambda E-A)^{-1})^{p+2}z=E((\lambda E-A)^{-1}E)^{p+1}(\lambda E-A)^{-1}z$
and thus
@@
$(E,A)$ $\sim(\tilde{R}E\tilde{P}^{-1},\tilde{R}A\tilde{P}^{-1})$ (19a)
$\sim\left(\begin{bmatrix}E_{1}&0\\ 0&E_{2}\end{bmatrix},\begin{bmatrix}A_{1}&0\\ 0&A_{2}\end{bmatrix}\right)$ (19b)
$\sim\left(\begin{bmatrix}I_{\mathcal{Z}^{1}}&0\\ 0&E_{2}A_{2}^{-1}\end{bmatrix},\begin{bmatrix}A_{1}E_{1}^{-1}&0\\ 0&I_{\mathcal{Z}^{2}}\end{bmatrix}\right).$ (19c)
The uniqueness of the Weierstraß form follows from the uniqueness of the nilpotency index [2, Prop. 6.2], which exists since the nilpotency index of $(E,A)$ is at most $p_{\rm rad}^{(E,A)}+1$ [2, Prop. 6.3]. ∎
@@
With (19c) one can rewrite the DAE $\frac{d}{dt}Ex=Ax$ as follows
$\frac{d}{dt}\begin{bmatrix}I_{\mathcal{Z}^{1}}&0\\ 0&E_{2}A_{2}^{-1}\end{bmatrix}\begin{bmatrix}z_{1}(t)\\ z_{2}(t)\end{bmatrix}=\begin{bmatrix}A_{1}E_{1}^{-1}&0\\ 0&I_{\mathcal{Z}^{2}}\end{bmatrix}\begin{bmatrix}z_{1}(t)\\ z_{2}(t)\end{bmatrix},$ (20)
on $\mathcal{Z}^{1}\times\mathcal{Z}^{2}$. Further, one can also show that
$(E,A)\sim\left(\begin{bmatrix}I_{\mathcal{X}^{1}}&0\\ 0&A_{2}^{-1}E_{2}\end{bmatrix},\begin{bmatrix}E_{1}^{-1}A_{1}&0\\ 0&I_{\mathcal{X}^{2}}\end{bmatrix}\right)$
and rewrite the DAE $E\frac{d}{dt}x=Ax$ in the following way
$\begin{bmatrix}I_{\mathcal{X}^{1}}&0\\ 0&A_{2}^{-1}E_{2}\end{bmatrix}\frac{d}{dt}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$ $=\begin{bmatrix}E_{1}^{-1}A_{1}&0\\ 0&I_{\mathcal{X}^{2}}\end{bmatrix}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix},$ (21)
on $\mathcal{X}^{1}\times\mathcal{X}^{2}$.
@@
Let $\mathcal{W}$ be a Hilbert space and $(A_{0},\mathcal{D}(A_{0}))$ generate a $C_{0}$-semigroup on $\mathcal{W}$. For $b\in\mathcal{D}(A_{0})$ and $c\in\mathcal{D}(A_{0}^{\ast})$ with $\langle b,c\rangle\neq 0$ define the operators $B_{u}=bu$, where $u\in\mathbb{C}$ and $Cz=\langle c,z\rangle$ for any $z\in\mathcal{W}$. We define the following DAE on $\mathcal{Z}=\mathcal{W}\times\mathbb{C}$ by